# Rourkela - 769008, India

### v1

@@ 3.3.1 Massless particles
Unlike the dyonic and Bardeen black holes studied in the previous subsections, since we are slowly increasing the pressure and taking it beyond the critical value, the Lyapunov exponent does not attain the same fixed constant value at large $T$. This is illustrated in Fig. 17(c). The reason is that changing the pressure (or cosmological constant) ultimately changes the AdS radius. If we take $l=1$ and work in the standard phase space, then $\lambda$ saturates to a constant value at large $T$ and $r_{h}$ in the Gauss-Bonnet black hole background.
@@ 3.3.2 Massive particles
3.4 AdS Black Holes in Massive Gravity
The study of AdS black holes with massive gravity is an exciting and dynamic field of theoretical research Fierz:1939ix ; deRham:2010ik ; deRham:2010kj ; Dvali:2000hr ; Bergshoeff:2009hq ; deRham:2014zqa ; Hinterbichler:2011tt . Giving the graviton a mass considerably changes general relativity by weakening it at large scales but nonetheless, it leads to the same predictions as general relativity at small scales. In principle, the massive gravity systems could explain the acceleration of the Universe without introducing the dark energy component, thereby making them highly popular in gravitation studies. By adding an effective mass to the graviton, the large graviton modifies the gravitational potential and affects how the black hole solutions behave. Compared to their general relativity counterparts, these black holes have unique characteristics such as altered horizon structures, thermodynamics, and dynamics. The study of massive gravity in AdS spaces has also revealed new information on thermodynamic phase transitions, the nature of gravitational waves, and the holographic dualities between massive gravity and their boundary theories. For more information and the current status of the massive gravity theories, we refer the readers to review papers deRham:2014zqa ; Hinterbichler:2011tt ; deRham:2016nuf . From the perspective of this paper, the AdS massive gravity also exhibits interesting phase transitions. Therefore, it is worthwhile to explore further the interplay between phase transitions and the Lyapunov exponent in these gravity systems.
Of the many avatars of massive gravity theories, we will examine a theory with Lorentz symmetry breaking. See Dubovsky:2004sg ; Rubakov:2008nh , for a review of Lorentz violating massive gravity theory. The metric of such massive gravity black holes was obtained in Bebronne:2009mz ; Comelli:2010bj , and their extended phase space thermodynamics were discussed in Fernando:2016sps .
@@ 3.4 AdS Black Holes in Massive Gravity
The above solution approaches the usual AdS-Schwarzschild black hole for large distances when $\omega>1$, whereas the Arnowitt-Deser-Misner mass of such solutions becomes divergent when $\omega<1$. Hence, here we will consider $\omega>1$. The parameter $\gamma$ can take two values $1$ or $-1$ Fernando:2016sps . In our analysis, we choose $\gamma=-1$, for which the geometry matches the RN-AdS charged black hole with two horizons.
@@
with the corresponding conjugate volume, $V$ given by
@@
Figure 22: The 3D plot of the effective potential $V_{\text{eff}}(r)$ as a function of horizon radius $\tilde{r_{h}}$ and orbit radius $r$ for the massive particle when $L=20$ , $P=0.001$, $\gamma=-1$, $Q=0.1$, and $\omega=3$. The brown and cyan curves projected in the lower part correspond to the unstable and stable equilibria of the time-like circular geodesics.
@@ 3.4.2 Massive particles
4 Critical exponent for charged black holes based on Lyapunov exponents
In the mean-field theories, critical exponents play an important role and are only defined as limiting power laws as $T\rightarrow T_{c}$ Goldenfeld2019Jun , which means that if any observable, say, $a(t=|T-T_{c}|)$ has critical exponent $\xi$ then,
@@ 4 Critical exponent for charged black holes based on Lyapunov exponents
Depending upon the physical observable under consideration, one can associate different mean-field critical exponents at the phase transition point, such as those associated with the specific heat, susceptibility, compressibility, etc. The one we are most interested in and whose existence forms the basis of the famous Landau theory is the order parameter critical exponent Wang2021May . The order parameter is the observable whose expectation value is nonzero below the transition temperature and zero above the transition temperature Aron/SciPostPhysLectNotes.11 . For the van der Waals liquid/gas phase transition, the density shows discontinuity between the two phases and is considered the order parameter. The order parameter critical exponent for it was found to be 1/2 van2004continuity .
The small/large black hole phase transition we saw in previous sections for different charged black holes is similar to the van der Waals liquid/gas transition Chamblin:1999hg . The temperature at which this first-order phase transition occurs is denoted by $T_{p}$ in our work. Let us also denote the Lyapunov exponent of small and large black holes, calculated at $T_{p}$, by $\lambda_{s}$ and $\lambda_{l}$, respectively and the one calculated at the second-order critical point by $\lambda_{c}$. Therefore, $\lambda_{s}=\lambda_{l}=\lambda_{c}$ at the second-order critical point. Following guo2022probing ; Wang2021May , one can consider the discontinuity in the Lyapunov exponent $\Delta\lambda=\lambda_{s}-\lambda_{l}$ at the first-order phase transition as an order parameter. In particular, we can try to find the critical exponent $\beta$, satisfying $\Delta\lambda$,
@@
4.1 Theoretical verification
We use the method illustrated in Banerjee:2012zm ; lyu2023probing to calculate this critical exponent $\beta$ related to the order parameter $\Delta\lambda$. First, we do a Taylor expansion of $\lambda(r_{h})$ up to the first order around the critical point as,
@@ 4.1 Theoretical verification
where the subscript $l$ and $s$ stands for large and small. Using Eqs. (81) and (4.1), we can get,
@@
Thus, we conclude that the critical exponent $\beta$ related to the order parameter $\Delta\lambda$ is $1/2$, the same as the van der Waals fluid van2004continuity . The same critical exponent of $1/2$ was also found in kumara2024lyapunov for another class of black holes, stressing its quite universal value.
@@ 5 Conclusions
Let us end this paper by pointing out a few limitations one faces by probing the black hole thermodynamic phase structure from Lyapunov exponents. An important point that we would like to mention, one which we think has not been discussed much, is that although the Lyapunov exponents of particles do seem to epitomize the presence of the black hole phase transition, they do not give any information about the transition temperature, especially if the transition is of the first order. The $\lambda~{}\text{vs}~{}T$ curves only emphasize the importance of $T_{1}$ around which $\lambda$ starts to show multi-valuedness. However, the first-order small/large black hole phase transition occurs at $T>T_{1}$, which cannot be determined from the thermal behaviour of the Lyapunov exponents. This should be contrasted with the information on the second-order critical points, which can be inferred from Lyapunov exponents. Another important point is that using the Lyapunov exponent presents a few issues, even within general relativity. In particular, since it amounts to the deviation of two nearby trajectories in time, the whole Lyapunov exponent analysis heavily depends on the considered time coordinate. This coordinate dependence can sometimes lead to wrong conclusions, for instance, by predicting zero Lyapunov exponents even for genuinely chaotic systems PhysRevLett.46.963 ; Barrow:1981sx ; Cornish:1996yg . However, this coordinate-dependent issue can be minimised when a preferred time direction exists, such as black hole geometries containing a time-like Killing vector, and meaningful results can be obtained Cornish:2003ig . In other words, if we are asymptotically far from the black hole, we can assign well-defined time coordinates in our observations. Our work also has a time-like killing vector for all the discussed black hole cases. Moreover, we compare all timescales in the same coordinate system, making the whole Lyapunov exponent analysis credible.
Acknowledgements.

### v2

@@ III.3.1 Massless particles
Unlike the dyonic and Bardeen black holes studied in the previous subsections, since we are slowly increasing the pressure and taking it beyond the critical value, the Lyapunov exponent does not attain the same fixed constant value at large $T$. This is illustrated in Fig. 17(c). The reason is that changing the pressure (or cosmological constant) ultimately changes the AdS radius. If we take $l=1$ and work in the standard phase space, then $\lambda$ saturates to a constant value at large $T$ and $r_{h}$ even in the Gauss-Bonnet black hole background as well.
@@ III.3.2 Massive particles
III.4 AdS Black Holes in Massive Gravity
The study of AdS black holes with massive gravity is an exciting and dynamic field of theoretical research [93, 94, 95, 96, 97, 98, 99]. Giving the graviton a mass considerably changes general relativity by weakening it at large scales, but nonetheless, it leads to the same predictions as general relativity at small scales. In principle, the massive gravity systems could explain the acceleration of the Universe without introducing the dark energy component, thereby making them highly popular in gravitation studies. By adding an effective mass to the graviton, the large graviton modifies the gravitational potential and affects how the black hole solutions behave. Compared to their general relativity counterparts, these black holes have unique characteristics such as altered horizon structures, thermodynamics, and dynamics. The study of massive gravity in AdS spaces has also revealed new information on thermodynamic phase transitions, the nature of gravitational waves, and the holographic dualities between massive gravity and their boundary theories. For more information and the current status of the massive gravity theories, we refer the readers to review papers [98, 99, 128]. From the perspective of this paper, the AdS massive gravity also exhibits interesting phase transitions. Therefore, it is worthwhile to explore further the interplay between phase transitions and the Lyapunov exponent in these gravity systems.
Of the many avatars of massive gravity theories, we will examine a theory with Lorentz symmetry breaking. See [129, 130], for a review of Lorentz violating massive gravity theory. The metric of such massive gravity black holes was obtained in [131, 132], and their extended phase space thermodynamics were discussed in [105].
@@ III.4 AdS Black Holes in Massive Gravity
The above solution approaches the usual AdS-Schwarzschild black hole for large distances when $\omega>1$, whereas the Arnowitt-Deser-Misner mass of such solutions becomes divergent when $\omega<1$. Hence, here we will consider $\omega>1$. The parameter $\gamma$ can take two values $1$ or $-1$ [105]. In our analysis, we choose $\gamma=-1$, for which the geometry matches the RN-AdS charged black hole with two horizons.
@@
with the corresponding conjugate volume $V$ given by
@@
Figure 22: The 3D plot of the effective potential $V_{\text{eff}}(r)$ as a function of horizon radius $\tilde{r_{h}}$ and orbit radius $r$ for the massive particle when $L=20$ , $P=0.001$, $\gamma=-1$, $Q=0.1$, and $\omega=3$. The brown and cyan curves projected in the lower part correspond to the unstable and stable equilibria of the timelike circular geodesics.
@@ III.4.2 Massive particles
IV Critical exponent for charged black holes based on Lyapunov exponents
In the mean-field theories, critical exponents play an important role and are only defined as limiting power laws as $T\rightarrow T_{c}$ [133], which means that if any observable, say, $a(t=|T-T_{c}|)$ has critical exponent $\xi$ then,
@@ IV Critical exponent for charged black holes based on Lyapunov exponents
Depending upon the physical observable under consideration, one can associate different mean-field critical exponents at the phase transition point, such as those associated with the specific heat, susceptibility, compressibility, etc. The one we are most interested in and whose existence forms the basis of the famous Landau theory is the order parameter critical exponent [134]. The order parameter is the observable whose expectation value is nonzero below the transition temperature and zero above the transition temperature [135]. For the van der Waals liquid/gas phase transition, the density shows discontinuity between the two phases and is considered the order parameter. The order parameter critical exponent for it was found to be 1/2 [136].
The small/large black hole phase transition we saw in previous sections for different charged black holes is similar to the van der Waals liquid/gas transition [137]. The temperature at which this first-order phase transition occurs is denoted by $T_{p}$ in our work. Let us also denote the Lyapunov exponent of small and large black holes, calculated at $T_{p}$, by $\lambda_{s}$ and $\lambda_{l}$, respectively and the one calculated at the second-order critical point by $\lambda_{c}$. Therefore, $\lambda_{s}=\lambda_{l}=\lambda_{c}$ at the second-order critical point. Following [82, 134], one can consider the discontinuity in the Lyapunov exponent $\Delta\lambda=\lambda_{s}-\lambda_{l}$ at the first-order phase transition as an order parameter. In particular, we can try to find the critical exponent $\beta$, satisfying $\Delta\lambda$,
@@
IV.1 Theoretical verification
We use the method illustrated in [138, 84] to calculate this critical exponent $\beta$ related to the order parameter $\Delta\lambda$. First, we do a Taylor expansion of $\lambda(r_{h})$ up to the first order around the critical point as,
@@ IV.1 Theoretical verification
where the subscript $l$ and $s$ stands for large and small. Using Eqs. (81) and (IV.1), we can get,
@@
Thus, we conclude that the critical exponent $\beta$ related to the order parameter $\Delta\lambda$ is $1/2$, the same as the van der Waals fluid [136]. The same critical exponent of $1/2$ was also found in [85] for another class of black holes, stressing its quite universal value.
@@ V Conclusions
Let us end this paper by pointing out a few limitations one faces by probing the black hole thermodynamic phase structure from Lyapunov exponents. An important point that we would like to mention, one which we think has not been discussed much, is that although the Lyapunov exponents of particles do seem to epitomise the presence of the black hole phase transition, they do not give any information about the transition temperature, especially if the transition is of the first order. The $\lambda~{}\text{vs}~{}T$ curves only emphasise the importance of $T_{1}$ around which $\lambda$ starts to show multi-valuedness. However, the first-order small/large black hole phase transition occurs at $T>T_{1}$, which cannot be determined from the thermal behaviour of the Lyapunov exponents. This should be contrasted with the information on the second-order critical points, which can be inferred from Lyapunov exponents. Another important point is that using the Lyapunov exponent presents a few issues, even within general relativity. In particular, since it amounts to the deviation of two nearby trajectories in time, the whole Lyapunov exponent analysis heavily depends on the considered time coordinate. This coordinate dependence can sometimes lead to wrong conclusions, for instance, by predicting zero Lyapunov exponents even for genuinely chaotic systems [139, 140, 141]. However, this coordinate-dependent issue can be minimised when a preferred time direction exists, such as black hole geometries containing a timelike Killing vector, and meaningful results can be obtained [107]. In other words, if we are asymptotically far from the black hole, we can assign well-defined time coordinates in our observations. Our work also has a timelike Killing vector for all the black hole cases discussed. Moreover, we compare all timescales in the same coordinate system, making the whole Lyapunov exponent analysis credible.
Acknowledgements